# Multi-Mode Array Filtering of Resonance Fluorescence

### v2

@@ IV.2 First-order correlation function
The normalized first-order correlation function for the multi-mode array filter in the steady-state limit is
@@
where we have introduced the non-normalized collective mode annihilation operator, $A=\sqrt{2N+1}\bar{A}$. Decomposing the filtered power spectrum into coherent and incoherent scattering components, the frequency-filtered incoherent power spectrum is then given by:
@@
$C_{1}$ $=\sum_{j=-N}^{N}\frac{\mathcal{E}_{j}^{*}}{2}\frac{\Delta_{-}+\Delta_{+}}{\gamma/2-\kappa+i\Delta\omega_{j}}$ (56a)
@@
While we do not give explicit expressions for the steady state expectation values in Eqs. (56-57), we can compute these numerically from the moment equations given in Appendix A.
@@ V.1.1 Varying the halfwidth
One gauge we can use to measure the efficacy of the multi-mode array filter is to investigate the initial value of the auto-correlation function. For very large filter bandwidths we expect an initial value close to zero, and for smaller bandwidths we expect to see different behavior from the secular approximation. In Fig. 10 we see the initial auto-correlation value tends towards zero as the filter halfwidth increases, while for smaller halfwidths it increases.
@@ V.1.2 Filtering within the peak
Figure 10 showed an increasing filter halfwidth resulted in auto-correlations that tended towards the bare atomic correlations. A decreasing halfwidth, however, resulted in correlations that deviated from the expected behavior of the secular approximation. In Fig. 12 we extend the range of filter halfwidths from $K=10^{-5}\gamma$ to $10^{2}\gamma$, highlighting the different regimes of behavior. Results for the single-mode filter (orange dashed line) have been reported in both theoretical and experimental work in the strong [48, 26] and weak driving regimes [49, 50, 51].
Figure 13: Intensity ratio of the single-mode (blue, solid) and multi-mode (orange, dashed) filtered coherent and incoherent spectra when resonant with the central peak, $\Delta\omega=0$, from Eq. (63). The inset is zoomed in on the region where $K\approx\Omega$, where there is a clear decrease in the incoherent intensity $I_{\mathrm{inc}}$, corresponding to the bunching peak in Fig. 12. Parameters are the same as in Fig. 10.
As the halfwidth decreases further, we see two different, yet related, regimes for the filtered central- and right-peaks. For vanishingly small halfwidths, $K\lll\gamma$ ($\sim 10^{-5}\gamma$), the input field is effectively a broadband field from the perspective of the filter. The input field then begins to resemble a thermal environment, resulting in the “thermalization” of the source field [18, 20, 24]. The side-peaks of the Mollow triplet are products of the incoherent scattering of the driving field by the atom. Hence as the filter halfwidth decreases, the initial value of the auto-correlation function tends towards that of thermal light, $g^{(2)}(\Omega,0;\Omega,0)=2$.
The central-peak differs, however, as there exists a coherent scattering component, which can be reasonably modeled by the Dirac-delta function. For vanishingly small halfwidths, more of the incoherent scattering is filtered out, eventually leaving only the infinitely narrow coherent component. The auto-correlation then tends towards second-order coherence, with $g^{(2)}(0,0;0,\tau)=1$.
@@
The multi-mode array filter produces similar results to the single-mode filter: large halfwidths give near perfect antibunching, intermediate halfwidths can isolate the target peaks, and vanishingly small halfwidths result in thermalization for the filtered right peak and second-order coherence for the filtered central peak. We do see some slight differences, such as a shift in the bunching peak for the filtered central peak, due to the improved frequency isolation. The most striking difference is the emergence of peaks at $K\sim\Omega$ for both the filtered central- and right-peaks, and a second peak at $K\sim 2\Omega$ for the filtered right-peak. Similar results can be seen in a perfect rectangular filter (see Figs. 4 and 5 of Ref. [35]). Kamide et al. suggest that these peaks are due to constructive interference in the cascaded decay channels.
@@
which we show for the filtered central peak in Fig. 13. When the halfwidth is approximately the same as the side-peak splitting, $K\approx\Omega$, there is a clear deviation in the single-mode and multi-mode filtered intensity ratios (highlighted in the inset of Fig. 13). As the frequency response of the multi-mode array filter expands and overlaps slightly with the side-peaks, there is a decrease in the incoherent scattering intensity, possibly due to the interference in the cascaded decay channels. We can infer a similar effect occurs when the filter is resonant with either side-peak. As the filter response first overlaps with the central peak, at $K\approx\Omega$, and then with the opposite side peak, at $K\approx 2\Omega$, there will be a similar decrease in the incoherent scattering. This, however, will need further investigation.
@@ V.2.1 Temporal evolution of peak-to-peak correlations
As previously discussed, the secular approximation in the atomic picture fails to account for the different time orderings of the de-excitation paths. The filtered correlations in Fig. 14 do not show the expected bunching for right-to-left filtered-correlations of Eq. (40a), or the second-order coherence of Eq. (40b). We instead see close agreement with the behavior expected from the equations derived by Schrama et al. [52], Eqs. (41).
We note here that we have chosen different halfwidths for the single- and multi-mode array filters, hence why the expected short-time behaviors differ. With a halfwidth roughly three times larger, however, the visible Rabi oscillations in the multi-mode-filtered cross-correlations are still significantly reduced when compared to the single-mode filter. Larger $K$ for the single-mode filter would result in much larger Rabi oscillations in the cross-correlations.
@@ V.2.2 Cross-correlations outside the Mollow triplet peaks
Figure 15: Initial value of the cross-correlation function, $g^{(2)}(\alpha,0;\beta,0)$, for the (a) single-mode filter, with $N=0$ and $K=\gamma$, and the (b) multi-mode array filters, with $N=80,K=5.5\gamma,\delta\omega=0.069\gamma$, and $\kappa=0.172\gamma$. The color code of the contour plot is: blue for $g^{(2)}(\alpha,0;\beta,0)<1$, white for $g^{(2)}(\alpha,0;\beta,0)=1$, and $g^{(2)}(\alpha,0;\beta,0)>1$. The driving amplitude is $\Omega=5\pi\gamma$
We are not limited to cross-correlations that are resonant with the three peaks of the Mollow triplet; we can choose the central resonance of filters $A$ and $B$ to be any frequency. Doing so allows us to uncover more interesting, and potentially useful, photon correlations. Figure 15 depicts landscapes of initial correlation values for the single- and multi-mode array filters, as the central frequencies of both filters are varied. We note that the results for the single-mode filter, Fig. 15(a) closely resembles the results from recent work [24, 26, 53, 25, 29], where the frequency filter was also modeled as a single-mode cavity.

### v3

@@ IV.2 First-order correlation function
The normalised first-order correlation function for the multi-mode array filter in the steady-state limit is
@@
where we have introduced the non-normalised collective mode annihilation operator, $A=\sqrt{2N+1}\bar{A}$. Decomposing the filtered power spectrum into coherent and incoherent scattering components, the frequency-filtered incoherent power spectrum is then given by:
@@
$C_{1}$ $=\sum_{j=-N}^{N}\frac{\mathcal{E}_{j}^{*}}{2}\frac{\Delta_{-}+\Delta_{+}}{\gamma/2-\kappa+i\Delta\omega_{j}},$ (55a)
@@
While we do not give explicit expressions for the steady state expectation values in Eqs. (55-56), we can compute these numerically from the moment equations given in Appendix A.
@@ V.1.1 Varying the halfwidth
One gauge we can use to measure the efficacy of the multi-mode array filter is to investigate the initial value of the auto-correlation function. For very large filter bandwidths we expect an initial value close to zero, and for smaller bandwidths we expect to see different behaviour from the secular approximation. In Fig. 10 we see the initial auto-correlation value tends towards zero as the filter halfwidth increases, while for smaller halfwidths it increases.
@@
We note here that the filter halfwidths in Fig. 11 were chosen via numerical optimisation where the Rabi oscillations were minimised, i.e, a close match to the secular approximations. In principle, one could also just pick a halfwidth for the multi-mode array filter that is half the frequency separation between the target peak and its nearest neighbour, e.g., $\Omega/2$ for the resonantly driven two-level atom, without substantially affecting the results.
@@ V.1.2 Filtering within the peak
Figure 10 showed an increasing filter halfwidth resulted in auto-correlations that tended towards the bare atomic correlations. A decreasing halfwidth, however, resulted in correlations that deviated from the expected behaviour of the secular approximation. In Fig. 12 we extend the range of filter halfwidths from $K=10^{-5}\gamma$ to $10^{2}\gamma$, highlighting the different regimes of behaviour. Results for the single-mode filter (orange dashed line) have been reported in both theoretical and experimental work in the strong [49, 27] and weak driving regimes [50, 51, 52].
@@
As the halfwidth decreases further, we see two different, yet related, regimes for the filtered central- and right-peaks. For vanishingly small halfwidths, $K\lll\gamma$ ($\sim 10^{-5}\gamma$), the input field is effectively a broadband field from the perspective of the filter. The input field then begins to resemble a thermal environment, resulting in the “thermalisation” of the source field [18, 20, 25]. The side-peaks of the Mollow triplet are products of the incoherent scattering of the driving field by the atom. Hence as the filter halfwidth decreases, the initial value of the auto-correlation function tends towards that of thermal light, $g^{(2)}(\Omega,0;\Omega,0)=2$.
The central-peak differs, however, as there exists a coherent scattering component, which can be reasonably modelled by the Dirac-delta function. For vanishingly small halfwidths, more of the incoherent scattering is filtered out, eventually leaving only the infinitely narrow coherent component. The auto-correlation then tends towards second-order coherence, with $g^{(2)}(0,0;0,\tau)=1$.
@@
The multi-mode array filter produces similar results to the single-mode filter: large halfwidths give near perfect antibunching, intermediate halfwidths can isolate the target peaks, and vanishingly small halfwidths result in thermalisation for the filtered right peak and second-order coherence for the filtered central peak. We do see some slight differences, such as a shift in the bunching peak for the filtered central peak, due to the improved frequency isolation. The most striking difference is the emergence of peaks at $K\sim\Omega$ for both the filtered central- and right-peaks, and a second peak at $K\sim 2\Omega$ for the filtered right-peak. Similar results can be seen in a perfect rectangular filter (see Figs. 4 and 5 of Ref. [36]). Kamide et al. suggest that these peaks are due to constructive interference in the cascaded decay channels.
@@
which we show for the filtered central peak in Fig. 13. When the halfwidth is approximately the same as the side-peak splitting, $K\approx\Omega$, there is a clear deviation in the single-mode and multi-mode filtered intensity ratios (highlighted in the inset of Fig. 13). As the frequency response of the multi-mode array filter expands and overlaps slightly with the sidepeaks, there is a decrease in the incoherent scattering intensity, possibly due to the interference in the cascaded decay channels. We can infer a similar effect occurs when the filter is resonant with either sidepeak. As the filter response first overlaps with the central peak, at $K\approx\Omega$, and then with the opposite side peak, at $K\approx 2\Omega$, there will be a similar decrease in the incoherent scattering. This, however, will need further investigation.
Figure 13: Intensity ratio of the multi-mode (blue, solid) and single-mode (orange, dashed) filtered coherent and incoherent spectra when resonant with the central peak, $\Delta\omega=0$, from Eq. (62). The inset is zoomed in on the region where $K\approx\Omega$, where there is a clear decrease in the incoherent intensity $I_{\mathrm{inc}}$, corresponding to the bunching peak in Fig. 12. Parameters are the same as in Fig. 10.
@@ V.2 Cross-correlations
Figure 15: Initial value of the cross-correlation function, $g^{(2)}(\alpha,0;\beta,0)$, for the (a) single-mode filter, with $N=0$ and $K=\gamma$, and the (b) multi-mode array filters, with $N=80,K=5.5\gamma,\delta\omega=0.069\gamma$, and $\kappa=0.172\gamma$. The colour code of the contour plot is: blue for $g^{(2)}(\alpha,0;\beta,0)<1$, white for $g^{(2)}(\alpha,0;\beta,0)=1$, and red for $g^{(2)}(\alpha,0;\beta,0)>1$. The driving amplitude is $\Omega=5\pi\gamma$
@@ V.2.1 Temporal evolution of peak-to-peak correlations
As previously discussed, the secular approximation in the atomic picture fails to account for the different time orderings of the de-excitation paths. The filtered correlations in Fig. 14 do not show the expected bunching for right-to-left filtered-correlations of Eq. (39a), or the second-order coherence of Eq. (39b). We instead see close agreement with the behaviour expected from the equations derived by Schrama et al. [53], Eqs. (40).
These theoretical results do slightly differ from experimental literature, as in Refs. [14, 43]. For the right-to-central peak cross-correlation, Fig. 14a, the central dip due to the destructive interference is seen in Figs. 12-14 of Ref. [43]. The experimental results, however, do not show any visible Rabi oscillations, possibly due to the weaker driving amplitude.
We note here that we have chosen different halfwidths for the single- and multi-mode array filters, with the same numerical optimisation method as in Fig. 11, hence why the expected short-time behaviours differ. With a halfwidth roughly three times larger, however, the visible Rabi oscillations in the multi-mode-filtered cross-correlations are still significantly reduced when compared to the single-mode filter. Larger $K$ for the single-mode filter would result in much larger Rabi oscillations in the cross-correlations.
@@ V.2.2 Cross-correlations outside the Mollow triplet peaks
We are not limited to cross-correlations that are resonant with the three peaks of the Mollow triplet; we can choose the central resonance of filters $A$ and $B$ to be any frequency. Doing so allows us to uncover more interesting, and potentially useful, photon correlations. Figure 15 depicts landscapes of initial correlation values for the single- and multi-mode array filters, as the central frequencies of both filters are varied. We note that the results for the single-mode filter, Fig. 15(a) closely resembles the results from recent work [25, 27, 54, 26, 30], where the frequency filter was also modelled as a single-mode cavity.
@@ VI Conclusion
In principle, this filtering system could be applied to any quantum optical source system. To calculate the complete set of filter-operator moment equations, the set of moment equations of the source system must be closed. We have also applied this method to a three-level ladder-type atom driven at two-photon resonance, with a manuscript currently in preparation [58].